OpenStudy (anonymous):

Hi. Can someone please help me with this physics problem? A faulty catapult launches a projectile straight up instead of forward. If its launch speed is 50 m/s, how high will the projectile be after 4 seconds? SHOW ALL WORK.

OpenStudy (anonymous):

I think this is how you do it: d=vit+(1/2)a(t^2) d=(50m/s)(4s)+(1/2)(9.8m/s/s)(4^2) d=200+78.4 d=278.4m Im not positive but I think this is it..

OpenStudy (anonymous):

thank you!